# GraphLoRA: Structure-Aware Contrastive Low-Rank Adaptation for Cross-Graph Transfer Learning

### v1

@@ B.2. Dataset Descriptions
GRACE 69.38±0.08 64.08±0.13 72.82±0.44 87.53±0.01 71.64±0.13
GRACE-P 69.36±0.17 52.68±0.66 64.38±0.08 85.76±0.02 73.45±0.12
@@ Appendix C Baselines
Appendix D Experimental settings
For supervised methods, we utilize labeled data from the target graph to train the GNN in a supervised manner. On the other hand, for graph contrastive learning methods and graph prompt learning methods, we first pre-train the GNN on the target graph in an unsupervised manner. Subsequently, we freeze the model parameters and fine-tune either a linear classifier or a graph prompt using the labels from the target graph. For transfer learning methods, we adopt a similar approach by pre-training the GNN on the source graph in an unsupervised manner. We then transfer the pre-trained model to the target graph and fine-tune a linear classifier using the labels from the target graph.
For all baselines, we utilize publicly available code provided by the authors. We employ the Adam (Kingma and Ba, 2015) optimizer for optimization, and the hyperparameters are tuned as suggested by the authors.
Appendix E Additional Experimental Results
E.1. Performance Comparison in the Public Setting and the 10-shot Setting
The comprehensive results of the comparison experiment in the public setting and the 10-shot setting are shown in Table 7.
E.2. Performance Comparison in the 5-shot Setting
The results of the comparison experiment in the 5-shot setting are shown in Table 8.
E.3. Ablation Studies in the 5-shot Setting
The results of the ablation experiment in the 5-shot setting are shown in Table 9.

### v2

@@ B.2. Dataset Descriptions
GPF 69.46±1.75 62.68±1.63 71.76±0.88 87.19±0.88 75.30±0.49
GraphPrompt 75.23±0.93 69.71±1.06 79.90±0.74 86.35±0.41 72.43±0.27
ProG 70.08±0.77 64.21±0.73 71.74±1.13 87.76±0.75 73.69±1.14
transfer GRACE-P 69.36±0.17 52.68±0.66 64.38±0.08 85.76±0.02 73.45±0.12
@@ Appendix C Baselines
  • GPF (Fang et al., 2023): GPF is a universal prompt-based tuning method for pre-trained GNN models, theoretically achieving the same effect as any form of prompting function.
  • GraphPrompt (Liu et al., 2023b): GraphPrompt introduces a unification framework by mapping different tasks to a common task template and proposes a learnable task-specific prompt vector to guide each downstream task in fully leveraging the pre-trained model.
  • ProG (Sun et al., 2023): ProG reformulates different-level tasks into unified ones and designs a multi-task prompting method for graph models.
@@
Appendix D Additional Experimental Results
D.1. Performance on Heterogeneous Graphs
We evaluate GraphLoRA on two heterogeneous graphs, Squirrel and Chameleon (Pei et al., 2020), with results presented in Table 9. The results show that GraphLoRA does not perform the best on heterogeneous graphs, which may be due to the Structure-aware Regularization module leveraging the homophily property of homogeneous graphs, making it less suitable for heterogeneous graphs. Nonetheless, GraphLoRA ranked 6th and 3rd among 14 methods on the two datasets, respectively, indicating respectable performance.
D.2. Performance with Different Pretraining Methods
We evaluate GraphLoRA with different pretraining methods, including CCA-SSG and HomoGCL, with results in Table 10. GraphLoRA performs best in most cases, achieving 13.66% better than CCA-SSG and 4.05% better than HomoGCL on average, demonstrating its effectiveness across various pretraining methods.
D.3. Performance in the 5-shot Setting
The results of the comparison experiment in the 5-shot setting are shown in Table 11. The results of the ablation experiment in the 5-shot setting are shown in Table 12.